OpenStudy (anonymous):

how do you solve the following system 3x^2+x-3y=-8 and x+3y=9 by elimination method step by step please

OpenStudy (anonymous):

Elimination method means that we want to remove 1 of the variables to be easier to solve In this case, I would like to remove the variable y Step 1: So let take addition of two equations for both sides \[3x^2+x-3y = -8\] \[x+3y = 9\] Take addition so we obtain a new equation \[3x^2 + 2x = 1\] Step 2: Solve this new equation (quadratic case) by any of known methods (factoring, or square, etc.) you can obtain x = -1 or x = 1/3 Step 3: Back substitution into x + 3y = 9 ==> y = 10/3 or y = 26/9

OpenStudy (anonymous):

how do you solve 3x2+2x=1 by using the quadratic equation thanks

OpenStudy (anonymous):

Did you study about a quadratic equation? There are several ways to solve a quadratic equation. I did it by using method of factoring. \[3x^2 + 2x - 1 = 0\] \[\rightarrow 3x^2 + 3x - x - 1 = 0\] \[\rightarrow 3x(x+1) - (x+1) = 0\] \[\rightarrow (3x-1)(x+1) = 0\] Until this step, you know what x is. More references to study of quadratic, you may be interested in mathisfun http://www.mathsisfun.com/algebra/quadratic-equation.html

OpenStudy (anonymous):

dont you have to multiply 3 to -1 before you can factor?

OpenStudy (anonymous):

I dont understand what do you mean by multiply 3 to -1. lol. Can you write it down?

OpenStudy (anonymous):